# Uniform irreducibility of Galois action on the $\ell$-primary part of Abelian $3$-folds of Picard type

### v1

@@ 1.1. The Bruhat-Tits tree of $\operatorname{U}_{3}$
The standard Iwahori subgroup of $G$ is defined as $I=K^{\circ}\cap K^{\prime}=\begin{pmatrix}\mathcal{O}^{\times}&\xi\mathcal{O}&\xi\mathcal{O}\\ \xi^{-1}\mathcal{P}&\mathcal{O}^{\times}&\mathcal{O}\\ \xi^{-1}\mathcal{P}&\mathcal{P}&\mathcal{O}^{\times}\end{pmatrix}\cap G$.
Finally let $I_{2,1}=K^{\circ}\cap\gamma K^{\circ}\gamma^{-1}=\begin{pmatrix}\mathcal{O}^{\times}&\xi\mathcal{O}&\xi\mathcal{O}\\ \xi^{-1}\mathcal{P}&\mathcal{O}^{\times}&\mathcal{O}\\ \xi^{-1}\mathcal{P}^{2}&\mathcal{P}&\mathcal{O}^{\times}\end{pmatrix}\cap G$, where $\gamma=\left(\begin{smallmatrix}\varpi^{-1}&&\\ &1&\\ &&\varpi\end{smallmatrix}\right)$.
The standard apartment in the Bruhat-Tits tree of $G$ is as follows:
@@
If $E$ is unramified, then $G\supset K^{\prime}\supset I\supset I_{2,1}\supset I_{2}$, where $I_{r}=\left(\begin{smallmatrix}\mathcal{O}^{\times}&\mathcal{O}&\mathcal{O}\\ p^{r}\mathcal{O}&\mathcal{O}^{\times}&\mathcal{O}\\ p^{r}\mathcal{O}&p^{r}\mathcal{O}&\mathcal{O}^{\times}\end{smallmatrix}\right)\cap G$.
@@ 1.2. Review of $L$-parameters and $A$-packets
The $L$-group (of the quasi-split form) of $\operatorname{U}(n)$ is given by $\operatorname{GL}_{n}(\mathbf{C})\rtimes W_{\mathbf{Q}_{p}}$ with the Weil group $W_{\mathbf{Q}_{p}}$ acting on $\operatorname{GL}_{n}(\mathbf{C})$ through its quotient $\operatorname{Gal}(E/\mathbf{Q}_{p})$ whose non-trivial element sends $g$ to $w_{n}{}^{t}g^{-1}w_{n}^{-1}$, where $w_{n}$ denotes the anti-diagonal matrix $(1,-1,1,\dots,(-1)^{n-1})$. By definition, an $L$-parameter for $\operatorname{U}(n)$ is a homomorphism $W_{\mathbf{Q}_{p}}\times\operatorname{SL}_{2}(\mathbf{C})\longrightarrow{}^{L}G$, but as one knows (see [13, §3]) it is equivalent to ask for its restriction
$\phi:W_{E}\times\operatorname{SL}_{2}(\mathbf{C})\longrightarrow\operatorname{GL}_{n}(\mathbf{C}),$
to be conjugate $(-1)^{n}$-dual, i.e., conjugate-orthogonal if $n$ is odd and conjugate-symplectic if $n$ is even. Recall that $\phi$ is conjugate-self-dual if $\bar{\phi}\simeq\phi^{\vee}$, or equivalently, if the induced representation $\operatorname{Ind}_{W_{E}}^{W_{\mathbf{Q}_{p}}}(\phi)$ is self-dual. Furthermore, $\phi$ is conjugate-orthogonal, resp. conjugate-symplectic, if it preserves a non-degenerate symmetric, resp. skew-symmetric, bilinear form. Note that while Schur’s Lemma implies that any irreducible self-dual (or conjugate-self-dual) parameter has a well defined sign, this need not be always the case for reducible parameters.
@@
The base change $\nu_{E}(z)=\nu(z/\overline{z})$ of a character $\nu$ of $E^{1}$ is conjugate-orthogonal and conversely any conjugate-orthogonal character of $E^{\times}$ is obtained in that way. For $\lambda$ a conjugate-symplectic character of $E^{\times}$, an example of key relevance to us is the conjugate-orthogonal representation
$(\lambda\otimes\operatorname{St})\oplus\nu_{E} :W_{E}\times\operatorname{SL}_{2}(\mathbf{C})\longrightarrow\operatorname{GL}_{3}(\mathbf{C}).$
It yields an $L$-parameter $\phi_{\lambda,\nu}$ of $G$, coming from an $L$-parameter of the (unique) cuspidal endoscopic subgroup $H=\operatorname{U}(1,1)\times\operatorname{U}(1)$ of $G$. The cardinality of the corresponding $L$-packet $\Pi_{L}(\phi_{\lambda,\nu})$ is given by the order of the centralizer in $^{L}G^{0}$ modulo the center which turns out to be $2$. More precisely, $\Pi_{L}(\phi_{\lambda,\nu})$ contains two discrete series representations $\pi_{2}$ and $\pi_{c}$ of $\operatorname{U}(3)$, exactly one of them, namely $\pi_{c}$, being supercuspidal (see [25, Chap. 12.2] where this $L$-packet is denoted $\Pi_{L}(\operatorname{St}_{H}(\xi))$). There is another endoscopic $L$-packet for $G$ consisting of a single non-tempered representation $\pi_{n}$ whose the $L$-parameter is given by
$\lambda|\cdot|_{E}^{1/2}\oplus\lambda|\cdot|_{E}^{-1/2}\oplus\nu_{E}:W_{E}\times\operatorname{SL}_{2}(\mathbf{C})\longrightarrow\operatorname{GL}_{3}(\mathbf{C}).$
Rogawski’s theory [25, 26] describes the automorphic representations contributing to the $\operatorname{H}^{1}$ of Shimura surfaces of Picard type in terms global Arthur packets (see [10, §3.1] for a summary). The corresponding local Arthur packet at $p$ has $2$ elements $\Pi(\lambda,\nu)=\{\pi_{n},\pi_{c}\}$ (see [25, §12.3.3], where $\pi_{c}$ is denoted $\pi^{s}$), and the restriction to $W_{E}$ of its $A$-parameter is given by
$(\lambda\otimes\mathbf{1}\otimes\operatorname{St})\oplus\nu_{E}:W_{E}\times\operatorname{SL}_{2}(\mathbf{C})\times\operatorname{SL}_{2}(\mathbf{C})\longrightarrow\operatorname{GL}_{3}(\mathbf{C}),$
@@
Crucial for us would be the description $\pi_{n}$ and $\pi_{2}$ as the Jordan–Hölder constituents of a principal series representation $\pi$. Indeed, by [26, §1], $\pi_{n}$ is the Langlands quotient of the (unitarily normalized) parabolic induction of the character
@@
with $\pi_{2}$ being the unique non-zero irreducible sub-representation. Moreover, the extension
@@
does not split, and the sub and quotient are switched when $\mu$ is replaced by $\mu^{w}(\bar{\alpha},\beta,\alpha^{-1})=\lambda(\bar{\alpha})\nu(\beta)|\alpha|_{E}^{-1/2}$. The Jacquet functor is exact and it sends $\pi_{2}$, resp. $\pi_{n}$, to $\mu\delta^{1/2}$, resp. $\mu^{w}\delta^{1/2}$, where $w$ is the non-trivial element of the Weyl group of $G$, and $\delta(\bar{\alpha},\beta,\alpha^{-1})=|\alpha|_{E}^{2}$ is the modulus character. Fixing a non-degenerate character of the unipotent radical $N$ of $B$, one knows by Rodier [24, Thm. 2] that the image of $\operatorname{Ind}_{B}^{G}(\mu)$ by a twisted version of the Jacquet functor, singling out generic representations, is a line. The later being also exact, this implies that exactly one amongst $\pi_{2}$ and $\pi_{n}$ is generic. Since $\pi_{n}$ is non-generic (see [25, p.174]), this implies that $\pi_{2}$ is generic.
@@ 1.3. The Gross subgroup $K^{\prime\prime}$
In this subsection, $E$ is ramified (hence $p$ is odd). Then $\mathcal{O}/\mathcal{P}=\mathbf{F}_{p}$ and $\mathcal{P}=\xi\cdot\mathcal{O}$. As $|\operatorname{PGL}_{2}(\mathbf{F}_{p})|=|\operatorname{SL}_{2}(\mathbf{F}_{p})|$ all vertices in the tree of $G$ have valence $p^{3}+1$ The map from $K^{\circ}$ to its reductive quotient $\overline{G}{}^{\circ}$ is obtained by reducing $\left(\begin{smallmatrix}\xi/2&&\\ &1&\\ &&1\end{smallmatrix}\right)^{-1}K^{\circ}\left(\begin{smallmatrix}\xi/2&&\\ &1&\\ &&1\end{smallmatrix}\right)$ modulo $\mathcal{P}$ and a direct computation shows that $\overline{G}{}^{\circ}$ is isomorphic to the orthogonal group $\mathrm{O}_{3}(\mathbf{F}_{p})$ with respect to the quadratic form represented by $\left(\begin{smallmatrix}&&1\\ &2&\\ 1&&\end{smallmatrix}\right)$.
Note that $\mathrm{O}_{3}(\mathbf{F}_{p})=\pm\mathrm{SO}_{3}(\mathbf{F}_{p})$ and the adjoint action on matrices $\left(\begin{smallmatrix}y&x\\ z&-y\end{smallmatrix}\right)$ preserving the determinant $(y^{2}+xz)$ allows us to identify $\operatorname{PGL}_{2}(\mathbf{F}_{p})$ and $\mathrm{SO}_{3}(\mathbf{F}_{p})$ as follows:
(4) $\begin{pmatrix}a&b\\ c&d\end{pmatrix}\mapsto\frac{1}{ad-bc}\begin{pmatrix}a^{2}&2ab&-b^{2}\\ ac&ad+bc&-bd\\ -c^{2}&-2cd&d^{2}\end{pmatrix}$.
It follows from that description that $\mathrm{SO}_{3}(\mathbf{F}_{p})$ is generated by the set
$\left\{\left(\begin{smallmatrix}&&-1\\ &-1&\\ -1&&\end{smallmatrix}\right),\left(\begin{smallmatrix}a&&\\ &1&\\ &&a^{-1}\end{smallmatrix}\right),\left(\begin{smallmatrix}1&0&0\\ c&1&0\\ -c^{2}&-2c&1\end{smallmatrix}\right)\Big|a\in\mathbf{F}_{p}^{\times},c\in\mathbf{F}_{p}\right\}.$
@@
Let $K^{\prime\prime}$ be the index $2$ subgroup of $K^{\circ}$ defined as the inverse image of the subgroup of $\mathrm{O}_{3}(\mathbf{F}_{p})$ generated by $\mathbf{1}$ and the image of $\operatorname{PSL}_{2}(\mathbf{F}_{p})$. Let $I^{\prime\prime}=K^{\prime\prime}\cap K^{\prime}\subset I$.
We recall that $E/\mathbf{Q}_{p}$ is a ramified quadratic extension with $p$ odd. A conjugate-symplectic character $\lambda$ of $E^{\times}$ is necessarily ramified and its restriction to $\mathbf{Z}_{p}^{\times}$ is given by its unique quadratic character. If $\lambda$ is tamely ramified, then its restriction to $\mathcal{O}^{\times}$ is also given by its unique quadratic character, and the equation $\lambda(\xi)^{2}=\lambda(-\xi\bar{\xi})=\lambda(-1)=(-1)^{(p-1)/2}$ shows that there are precisely two such characters.
@@
As $\mu_{|(T\cap I^{\prime\prime})}=\mu^{w}_{|(T\cap I^{\prime\prime})}=\mathbf{1}$, applying the Jacquet functor to the exact sequence of admissible $G$-representations (2) allows one to see that both $\pi_{2}$ and $\pi_{n}$ have non-trivial $I^{\prime\prime}$-invariants. Moreover, as $\big|B\backslash G/I^{\prime\prime}\big|=\big|(B\cap K^{\prime\prime})\backslash K^{\prime\prime}/I^{\prime\prime}\big|=2$, both $\pi_{2}^{I^{\prime\prime}}$ and $\pi_{n}^{I^{\prime\prime}}$ must be $1$-dimensional. By Iwasawa decomposition, the restriction of $\operatorname{Ind}_{B}^{G}(\mu)$ to $K^{\prime\prime}$ is given by $\operatorname{Ind}_{B\cap K^{\prime\prime}}^{K^{\prime\prime}}(\mu)$, hence the line $\operatorname{Ind}_{B}^{G}(\mu)^{K^{\prime\prime}}$ admits a basis $f$ uniquely characterized by $f_{|K^{\prime\prime}}=\mathbf{1}_{K^{\prime\prime}}$. It follows that $\dim\pi_{n}^{K^{\prime\prime}}+\dim\pi_{2}^{K^{\prime\prime}}=1$ and we will show that $\pi_{2}^{K^{\prime\prime}}=\{0\}$.
@@
$\langle v,f^{\vee}\rangle=\frac{1}{\sqrt{\operatorname{vol}(K^{\prime\prime})}}\int_{K^{\prime\prime}}v(k)dk.$
@@
$\operatorname{vol}(K^{\prime\prime}\gamma^{n}K^{\prime\prime})\big|\langle\gamma^{n}\cdot f,f^{\vee}\rangle\big|^{2}=[K^{\prime\prime}:(K^{\prime\prime}\cap\gamma^{n}K^{\prime\prime}\gamma^{-n})]\Big|\int_{K^{\prime\prime}}f(k\gamma^{n})dk\Big|^{2}.$
@@
$\Phi_{n}=p^{5n/2}\cdot\Big|\int_{K^{\prime\prime}}f(\gamma^{-n}k\gamma^{n})dk\Big|.$
@@
For all $n\geqslant 1$ one has $\big|\int_{K^{\prime\prime}\backslash K^{\prime\prime}_{2n}}f(\gamma^{-n}k\gamma^{n})dk\big|\leqslant\operatorname{vol}(I^{\prime\prime})\cdot(\sqrt{p}+1)p^{-2n}$ and $\big|\int_{K^{\prime\prime}_{2n}}f(\gamma^{-n}k\gamma^{n})dk\big|=\operatorname{vol}(I^{\prime\prime})\cdot p^{1-2n}$.
@@
The last row of an element $k\in K^{\prime\prime}$ is given by $(0,0,1)\cdot k=(\xi^{-1}\cdot c_{1}(k),c_{2}(k),c_{3}(k))$ with $(c_{1}(k),c_{2}(k),c_{3}(k))\in\mathcal{O}^{3}\backslash(\xi\mathcal{O})^{3}$. For $j\geqslant 0$ we let
$K^{\prime\prime}_{j}=\left\{k\in K^{\prime\prime}\big|c_{1}(k)\in\xi^{j}\mathcal{O}\right\}\text{ and }K^{\prime\prime\times}_{j}=K^{\prime\prime}_{j}\backslash K^{\prime\prime}_{j+1}.$
Note that $K^{\prime\prime}_{0}=K^{\prime\prime}$ and $K^{\prime\prime}_{1}=I^{\prime\prime}$. We use the partition $K^{\prime\prime}\backslash K^{\prime\prime}_{2n}=K_{0}^{\prime\prime\times}\coprod K^{\prime\prime\times}_{1}\coprod\dots\coprod K^{\prime\prime\times}_{2n-1}$ to compute the first integral and $K^{\prime\prime}_{2n}=K^{\prime\prime\times}_{2n}\coprod K^{\prime\prime}_{2n+1}$ for the second.
@@
$\left|\int_{K^{\prime\prime}\backslash K^{\prime\prime}_{2n}}f(\gamma^{-n}k\gamma^{n})dk\right|\leqslant\operatorname{vol}(K^{\prime\prime}\backslash I^{\prime\prime})\cdot p^{-3n}+\operatorname{vol}(I^{\prime\prime})\sum_{j=1}^{2n-1}p^{\frac{1}{2}j-3n}(p-1)=$
@@
Since $\operatorname{vol}(I^{\prime\prime})\cdot p^{1-2n}=\operatorname{vol}(K^{\prime\prime}_{2n})$, in order to complete the proof of the lemma, it suffices to show that $f(\gamma^{-n}k\gamma^{n})$ in constant on $k\in K^{\prime\prime}_{2n}$. This is evident for $k\in K^{\prime\prime}_{2n+1}$, as then $k$ and $\gamma^{-n}k\gamma^{n}$ both belong to $I^{\prime\prime}$ and share same determinant and lower right coefficient $c_{3}$, implying that $f(\gamma^{-n}k\gamma^{n})=f(k)$. Miraculously, as one can see from (4), this remains true for $k\in K^{\prime\prime}_{2n}\backslash K^{\prime\prime}_{2n+1}$ as well, i.e. even thought $\I\gamma^{-n}k\gamma^{n}\in K^{\circ}\backslash$ I, the fact that $c_{3}(\gamma^{-n}k\gamma^{n})=c_{3}(k)$ still implies that $\gamma^{-n}k\gamma^{n}\in K^{\prime\prime}$. ∎

### v2

@@ 1.1. The Bruhat-Tits tree of $\operatorname{U}(3)$
The standard Iwahori subgroup of $G$, defined as $I=K^{\circ}\cap K^{\prime}=\left(\begin{smallmatrix}\mathcal{O}^{\times}&\mathcal{O}&\mathcal{O}\\ \mathcal{P}&\mathcal{O}^{\times}&\mathcal{O}\\ \mathcal{P}&\mathcal{P}&\mathcal{O}^{\times}\end{smallmatrix}\right)\cap G$, is the stabilizer of a chamber in the standard apartment in the Bruhat-Tits tree of $G$:
@@
where $\gamma=\left(\begin{smallmatrix}\varpi^{-1}&&\\ &1&\\ &&\varpi\end{smallmatrix}\right)$ and $I_{2,1}=K^{\circ}\cap\gamma K^{\circ}\gamma^{-1}=\left(\begin{smallmatrix}\mathcal{O}^{\times}&\mathcal{O}&\mathcal{O}\\ \mathcal{P}&\mathcal{O}^{\times}&\mathcal{O}\\ \mathcal{P}^{2}&\mathcal{P}&\mathcal{O}^{\times}\end{smallmatrix}\right)\cap G$.
One has $G\supset K^{\prime}\supset I\supset I_{2,1}\supset I_{2}$, where $I_{r}=\left(\begin{smallmatrix}\mathcal{O}^{\times}&\mathcal{O}&\mathcal{O}\\ \mathcal{P}^{r}&\mathcal{O}^{\times}&\mathcal{O}\\ \mathcal{P}^{r}&\mathcal{P}^{r}&\mathcal{O}^{\times}\end{smallmatrix}\right)\cap G$. Finally, we let $K_{1}\subset K^{\circ}$ be the principal congruence subgroup of matrices $\equiv\mathbf{1}_{3}\pmod{\mathcal{P}}$ and $K_{T}=\left(\begin{smallmatrix}\mathcal{O}^{\times}&\mathcal{P}&\mathcal{P}\\ \mathcal{P}&\mathcal{O}^{\times}&\mathcal{P}\\ \mathcal{P}&\mathcal{P}&\mathcal{O}^{\times}\end{smallmatrix}\right)\cap G$.
@@ 1.2. Review of $L$-parameters and $A$-packets
The $L$-group of $\operatorname{U}(n)$ is given by $\operatorname{GL}(n,\mathbf{C})\rtimes W_{\mathbf{Q}_{p}}$ with the Weil group $W_{\mathbf{Q}_{p}}$ acting on $\operatorname{GL}(n,\mathbf{C})$ through its quotient $\operatorname{Gal}(E/\mathbf{Q}_{p})$ whose non-trivial element sends $g$ to $w_{n}{}^{t}g^{-1}w_{n}^{-1}$, where $w_{n}$ denotes the anti-diagonal matrix $(1,-1,1,\dots,(-1)^{n-1})$. By definition, an $L$-parameter for the quasi-split $\operatorname{U}(n)$ is a homomorphism $W_{\mathbf{Q}_{p}}\times\operatorname{SL}(2,\mathbf{C})\longrightarrow\operatorname{GL}(n,\mathbf{C})\rtimes W_{\mathbf{Q}_{p}}$, but as one knows (see [14, §3]) it is equivalent to ask for its restriction
$\phi:W_{E}\times\operatorname{SL}(2,\mathbf{C})\longrightarrow\operatorname{GL}(n,\mathbf{C}),$
to be conjugate-orthogonal if $n$ is odd and conjugate-symplectic if $n$ is even. Recall that $\phi$ is conjugate-self-dual if $\overline{\phi}\simeq\phi^{\vee}$, or equivalently, if the induced representation $\operatorname{Ind}_{W_{E}}^{W_{\mathbf{Q}_{p}}}(\phi)$ is self-dual. Furthermore, $\phi$ is conjugate-orthogonal, resp. conjugate-symplectic, if it preserves a non-degenerate symmetric, resp. skew-symmetric, bilinear form. Note that while Schur’s Lemma implies that any irreducible self-dual (or conjugate-self-dual) parameter has a well defined sign, this need not be always the case for reducible parameters.
@@
The base change $\nu_{E}(z)=\nu(z/\overline{z})$ from $\operatorname{U}(1)$ to $\operatorname{GL}(1)/E$ of a character $\nu$ of $E^{1}$ is conjugate-orthogonal and conversely any conjugate-orthogonal character of $E^{\times}$ is obtained in that way. For $\lambda$ a conjugate-symplectic character of $E^{\times}$, the conjugate-orthogonal representation
$(\lambda\otimes\operatorname{St})\oplus\nu_{E} :W_{E}\times\operatorname{SL}(2,\mathbf{C})\longrightarrow\operatorname{GL}(3,\mathbf{C})$
would be of key relevance to us. It yields an $L$-parameter $\phi_{\lambda,\nu}$ of $G$, coming from an $L$-parameter of the (unique) cuspidal endoscopic subgroup $H=\operatorname{U}(1,1)\times\operatorname{U}(1)$ of $G$. The cardinality of the corresponding $L$-packet $\Pi_{L}(\phi_{\lambda,\nu})$ is given by the order of the centralizer of $\phi_{\lambda,\nu}$ (modulo center) which turns out to be $2$. More precisely, $\Pi_{L}(\phi_{\lambda,\nu})$ contains two discrete series representations $\pi_{2}$ and $\pi_{c}$ of $\operatorname{U}(3)$, exactly one of them, namely $\pi_{c}$, being supercuspidal (see [26, Chap. 12.2] where this $L$-packet is denoted $\Pi_{L}(\operatorname{St}_{H}(\xi))$). There is another endoscopic $L$-packet for $G$ consisting of a single non-tempered representation $\pi_{n}$ whose the $L$-parameter is given by
$\lambda|\cdot|_{E}^{1/2}\oplus\lambda|\cdot|_{E}^{-1/2}\oplus\nu_{E}:W_{E}\times\operatorname{SL}(2,\mathbf{C})\longrightarrow\operatorname{GL}(3,\mathbf{C}).$
Rogawski’s theory [26, 27] describes the automorphic representations contributing to the $\operatorname{H}^{1}$ of Shimura surfaces of Picard type in terms global Arthur packets (see [11, §3.1] for a summary). The corresponding local Arthur packet at $p$ has $2$ elements $\Pi(\lambda,\nu)=\{\pi_{n},\pi_{c}\}$ (see [26, §12.3.3], where $\pi_{c}$ is denoted $\pi^{s}$), and the restriction to $W_{E}$ of its $A$-parameter is given by
$(\lambda\otimes\mathbf{1}\otimes\operatorname{St})\oplus\nu_{E}:W_{E}\times\operatorname{SL}(2,\mathbf{C})\times\operatorname{SL}(2,\mathbf{C})\longrightarrow\operatorname{GL}(3,\mathbf{C}),$
@@
Crucial for us would be the description $\pi_{n}$ and $\pi_{2}$ as the Jordan–Hölder constituents of a principal series representation $\pi$. Indeed, by [27, §1], $\pi_{n}$ is the Langlands quotient of the (unitarily normalized) parabolic induction of the character
@@
with $\pi_{2}$ the unique non-zero irreducible sub-representation. The sub and quotient are switched when $\mu$ is replaced by $\mu^{\mathsf{w}}(\bar{\alpha},\beta,\alpha^{-1})=\lambda(\bar{\alpha})\nu(\beta)|\alpha|_{E}^{-1/2}$, where $\mathsf{w}=\left(\begin{smallmatrix}&&1\\ &1&\\ 1&&\end{smallmatrix}\right)$ is the non-trivial element of the Weyl group of $G$. The Jacquet functor $\pi\to\pi_{N}$ is exact and it sends $\pi_{2}$ (resp. $\pi_{n}$) to $\mu\delta^{1/2}$ (resp. $\mu^{\mathsf{w}}\delta^{1/2}$), where $\delta(\bar{\alpha},\beta,\alpha^{-1})=|\alpha|_{E}^{2}$ is the modulus character. The extension
@@
does not split since, by Frobenuis reciprocity, one has
$\operatorname{End}_{G}\left(\operatorname{Ind}_{B}^{G}(\mu)\right)=\operatorname{Hom}_{B}\left(\operatorname{Ind}_{B}^{G}(\mu),\mu\delta^{1/2}\right)=\operatorname{Hom}_{T}\left(\operatorname{Ind}_{B}^{G}(\mu)_{N},\mu\delta^{1/2}\right)\simeq\mathbf{C}.$
One knows by Rodier [25, Thm. 2] that the image of $\operatorname{Ind}_{B}^{G}(\mu)$ by a twisted (by a non-degenerate character of $N$) Jacquet functor, singling out generic representations, is a line. Since $\pi_{n}$ is non-generic (see [26, p.174]), the exactness of this functor implies that $\pi_{2}$ is generic.
@@ 1.3. The Gross subgroup $K^{\prime\prime}$
In this subsection, $E$ is assumed ramified (hence $p$ is odd), $\mathcal{O}/\mathcal{P}=\mathbf{F}_{p}$ and $\mathcal{P}=(\xi)$. As $|\operatorname{PGL}(2,\mathbf{F}_{p})|=|\operatorname{SL}(2,\mathbf{F}_{p})|$ all vertices in the tree of $G$ have valence $p^{3}+1$. The reductive quotient $\overline{G}{}^{\circ}$ of $K^{\circ}$ is isomorphic to the orthogonal group $\mathrm{O}(3,\mathbf{F}_{p})$ with respect to the quadratic form represented by $\left(\begin{smallmatrix}&&1\\ &1&\\ 1&&\end{smallmatrix}\right)$. The adjoint action on matrices $\left(\begin{smallmatrix}y&x\\ z&-y\end{smallmatrix}\right)$ preserving the determinant $(y^{2}+xz)$ allows us to identify $\operatorname{PGL}(2,\mathbf{F}_{p})$ and $\mathrm{SO}(3,\mathbf{F}_{p})$ as follows:
(4) $\begin{pmatrix}a&b\\ c&d\end{pmatrix}\mapsto\frac{1}{ad-bc}\begin{pmatrix}a^{2}&-ab&-b^{2}/2\\ -2ac&ad+bc&bd\\ -2c^{2}&2cd&d^{2}\end{pmatrix}$.
One has $\mathrm{O}(3,\mathbf{F}_{p})=\pm\mathbf{1}_{3}\cdot\mathrm{SO}(3,\mathbf{F}_{p})$. By the above description, $\mathrm{SO}(3,\mathbf{F}_{p})$ is generated by the set
$\left\{\left(\begin{smallmatrix}&&-1/2\\ &-1&\\ -2&&\end{smallmatrix}\right),\left(\begin{smallmatrix}a&&\\ &1&\\ &&a^{-1}\end{smallmatrix}\right),\left(\begin{smallmatrix}1&-b&-b^{2}/2\\ &1&b\\ &&1\end{smallmatrix}\right)\Big|a\in\mathbf{F}_{p}^{\times},b\in\mathbf{F}_{p}\right\}.$
@@
Let $K^{\prime\prime}$ be the index $2$ subgroup of $K^{\circ}$ defined as the inverse image of the subgroup of $\mathrm{O}(3,\mathbf{F}_{p})$ generated by $\mathbf{1}_{3}$ and the image of $\operatorname{PSL}(2,\mathbf{F}_{p})$. Let $I^{\prime\prime}=K^{\prime\prime}\cap K^{\prime}\subset I$.
As $E/\mathbf{Q}_{p}$ is a ramified quadratic extension with $p$ odd, a conjugate-symplectic character $\lambda$ of $E^{\times}$ is necessarily ramified and its restriction to $\mathbf{Z}_{p}^{\times}$ is given by its unique quadratic character. If $\lambda$ is tamely ramified, then its restriction to $\mathcal{O}^{\times}$ is also given by its unique quadratic character, and the equation $\lambda(\xi)^{2}=\lambda(-\xi\bar{\xi})=\lambda(-1)=(-1)^{(p-1)/2}$ shows that there are precisely two such characters.
@@
As $\mu_{|(T\cap I^{\prime\prime})}=\mu^{\mathsf{w}}_{|(T\cap I^{\prime\prime})}=\mathbf{1}$, applying the Jacquet functor to the exact sequence of admissible $G$-representations (2) allows one to see that both $\pi_{2}$ and $\pi_{n}$ have non-trivial $I^{\prime\prime}$-invariants. Moreover, as $\big|B\backslash G/I^{\prime\prime}\big|=\big|(B\cap K^{\prime\prime})\backslash K^{\prime\prime}/I^{\prime\prime}\big|=2$, both $\pi_{2}^{I^{\prime\prime}}$ and $\pi_{n}^{I^{\prime\prime}}$ must be $1$-dimensional. By Iwasawa decomposition, the restriction of $\operatorname{Ind}_{B}^{G}(\mu)$ to $K^{\prime\prime}$ is given by $\operatorname{Ind}_{B\cap K^{\prime\prime}}^{K^{\prime\prime}}(\mu)$, hence the line $\operatorname{Ind}_{B}^{G}(\mu)^{K^{\prime\prime}}$ admits a basis $f$ uniquely characterized by $f_{|K^{\prime\prime}}=\mathbf{1}_{K^{\prime\prime}}$. It follows that $\dim\pi_{n}^{K^{\prime\prime}}+\dim\pi_{2}^{K^{\prime\prime}}=1$ and we will show that $\pi_{2}^{K^{\prime\prime}}=\{0\}$.
@@
$\langle v,f^{\vee}\rangle=\frac{1}{\sqrt{\operatorname{vol}(K^{\prime\prime})}}\int\limits_{K^{\prime\prime}}v(k)dk.$
@@
$\operatorname{vol}(K^{\prime\prime}\gamma^{n}K^{\prime\prime})\big|\langle\gamma^{n}\cdot f,f^{\vee}\rangle\big|^{2}=[K^{\prime\prime}:(K^{\prime\prime}\cap\gamma^{n}K^{\prime\prime}\gamma^{-n})]\Big|\int\limits_{K^{\prime\prime}}f(k\gamma^{n})dk\Big|^{2}.$
@@
$\Phi_{n}=p^{5n/2}\cdot\Big|\int\limits_{K^{\prime\prime}}f(\gamma^{-n}k\gamma^{n})dk\Big|.$
@@
For all $n\geqslant 1$ one has $\left|\int\limits_{K^{\prime\prime}{\smallsetminus}K^{\prime\prime}_{2n}}f(\gamma^{-n}k\gamma^{n})dk\right|\leqslant\operatorname{vol}(I^{\prime\prime})\cdot(\sqrt{p}+1)p^{-2n}$ and $\left|\int\limits_{K^{\prime\prime}_{2n}}f(\gamma^{-n}k\gamma^{n})dk\right|=\operatorname{vol}(I^{\prime\prime})\cdot p^{1-2n}$.
@@
The last row of an element $k\in K^{\prime\prime}$ is given by $(0,0,1)\cdot k=(c_{1}(k),c_{2}(k),c_{3}(k))\in\mathcal{O}^{3}{\smallsetminus}\mathcal{P}^{3}$. For $j\geqslant 0$ we let
$K^{\prime\prime}_{j}=\left\{k\in K^{\prime\prime}\big|c_{1}(k)\in\mathcal{P}^{j}\right\}\text{ and }K^{\prime\prime\times}_{j}=K^{\prime\prime}_{j}{\smallsetminus}K^{\prime\prime}_{j+1}.$
Note that $K^{\prime\prime}_{0}=K^{\prime\prime}$ and $K^{\prime\prime}_{1}=I^{\prime\prime}$. We use the partition $K^{\prime\prime}{\smallsetminus}K^{\prime\prime}_{2n}=K_{0}^{\prime\prime\times}\coprod K^{\prime\prime\times}_{1}\coprod\dots\coprod K^{\prime\prime\times}_{2n-1}$ to compute the first integral and $K^{\prime\prime}_{2n}=K^{\prime\prime\times}_{2n}\coprod K^{\prime\prime}_{2n+1}$ for the second.
@@
$\left|\int\limits_{K^{\prime\prime}{\smallsetminus}K^{\prime\prime}_{2n}}f(\gamma^{-n}k\gamma^{n})dk\right|\leqslant\operatorname{vol}(K^{\prime\prime}{\smallsetminus}I^{\prime\prime})\cdot p^{-3n}+\operatorname{vol}(I^{\prime\prime})\sum_{j=1}^{2n-1}p^{\frac{1}{2}j-3n}(p-1)=$
@@
Since $\operatorname{vol}(I^{\prime\prime})\cdot p^{1-2n}=\operatorname{vol}(K^{\prime\prime}_{2n})$, in order to complete the proof of the lemma, it suffices to show that $f(\gamma^{-n}k\gamma^{n})$ is constant on $k\in K^{\prime\prime}_{2n}$. This is evident for $k\in K^{\prime\prime}_{2n+1}$, as then $k$ and $\gamma^{-n}k\gamma^{n}$ both belong to $I^{\prime\prime}$ and share same determinant and lower right coefficient $c_{3}$, implying that $f(\gamma^{-n}k\gamma^{n})=f(k)$. Miraculously, as one can see from (4), this remains true for $k\in K^{\prime\prime}_{2n}{\smallsetminus}K^{\prime\prime}_{2n+1}$ as well, i.e. even though $\gamma^{-n}k\gamma^{n}\in K^{\circ}{\smallsetminus}I$, the fact that $c_{3}(\gamma^{-n}k\gamma^{n})=c_{3}(k)$ still implies that $\gamma^{-n}k\gamma^{n}\in K^{\prime\prime}$. ∎